OpenStudy (anonymous):

Prove: sin θ - sin θ�cos2 θ = sin3 θ. You must show all work. Please help

OpenStudy (anonymous):

@am!rah

OpenStudy (aum):

Not clear what is between sin θ (?????) cos2 θ

OpenStudy (anonymous):

i have no clue it was in the question

OpenStudy (aum):

Can you post a screenshot of the question?

OpenStudy (turingtest):

the ¿ symbol was in the question?

OpenStudy (anonymous):

sure 1 minute

OpenStudy (anonymous):

there ya go

OpenStudy (turingtest):

ok that is a typo, end of story. contact whoever you can about that question because it is meaningless we might be able to guess what they meant if we think about it though...

OpenStudy (anonymous):

alright can you help me with one more

OpenStudy (turingtest):

sure

OpenStudy (anonymous):

how would you transform f(x) = 4 sin(2x - π) + 3 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (turingtest):

use the fact that\[\sin u=\cos(\frac\pi2-u)\]